OpenStudy (anonymous):

please write in scientific notation (2 x 10*7*)(1.3 x 10-*3*) Inside the *'s are exponents

OpenStudy (anonymous):

be a good boy

OpenStudy (anonymous):

So\[(2\times10^{7})(1.3\times10^{-3})\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

just pop that sucker into a calculator.

OpenStudy (anonymous):

what do i do with exponents i know its 2.6x10?

OpenStudy (anonymous):

i dont know the expontent answer

OpenStudy (unklerhaukus):

\[(2\times10^{7})(1.3\times10^{-3})\]\[=(2\times1.3)(10^7\times10^{-3})\]\[=2.6\times10^4\]

OpenStudy (anonymous):

so i subtract the expontent?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

wait so how would you do this (4 x 102)(15 x 103)

OpenStudy (unklerhaukus):

no add exponents

OpenStudy (anonymous):

2 and 3 are exponents

OpenStudy (anonymous):

the oens after the 10's

OpenStudy (unklerhaukus):

(7)+(-3)=4

OpenStudy (anonymous):

\[(2* 10^7)(1.3 * 10^{-3})=2*1.3*10^7*10^{-3}=2.6*10^4\]Clearly, not typing fast enough.....

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but can you help with this one

OpenStudy (anonymous):

so i have two examples?

OpenStudy (unklerhaukus):

\[(4 \times 10^2)(15 \times 10^3)\] \[=(4 \times15)( 10^3\times 10^2)\] \[=(4 \times15)( 10^{3+2})\]

OpenStudy (anonymous):

oh thanks so much

OpenStudy (anonymous):

so 60x10*5*

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

yeap, it is that easy

OpenStudy (anonymous):

well heres the prob

OpenStudy (unklerhaukus):

but i would write that as 60x10^5

OpenStudy (anonymous):

the answer says 60x10*6*

OpenStudy (anonymous):

you sure you dont multiply the exponents?

OpenStudy (unklerhaukus):

\[60\times 10^5=6\times10^6\]

OpenStudy (anonymous):

that is a choice too

OpenStudy (unklerhaukus):

\[x^n\times x^m=x^{n+m}\]

OpenStudy (anonymous):

Simplify: (x - 8)2 please help

OpenStudy (anonymous):

?????

OpenStudy (unklerhaukus):

ok for scientific notation one numeral before the decimal place is correct

OpenStudy (anonymous):

@UnkleRhaukus u my hero luv yah boy or girl no weird

OpenStudy (unklerhaukus):

(x - 8)^2 \[=(x-8)^2=(x-8)(x-8)=x(x-8)-8(x-8)=\cdots\]

OpenStudy (anonymous):

uncle

OpenStudy (anonymous):

help me with this then please?

OpenStudy (unklerhaukus):

; )

OpenStudy (anonymous):

its a division one

OpenStudy (anonymous):

im a little girl and thank u

OpenStudy (unklerhaukus):

did you complete that last one @davidbecerra ?

OpenStudy (anonymous):

yes answer was 6x10*6*

OpenStudy (anonymous):

one day @UnkleRhaukus we will be married and rich luv yah again

OpenStudy (anonymous):

help with this please

OpenStudy (unklerhaukus):

one day,

OpenStudy (unklerhaukus):

yer

OpenStudy (anonymous):

im single and adult text me plz

OpenStudy (anonymous):